OpenStudy (anonymous):

please help ;/ How long, to the nearest year, will it take me to become a millionaire if I invest $3000 at 12% interest compounded continuously? HINT [See Example 3.]

OpenStudy (jdoe0001):

have you covered logarithms yet?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but I'm having difficulty

OpenStudy (anonymous):

i thought it was 96 yr but nothing

OpenStudy (jdoe0001):

well, you're given the principal and the rate you'd just need to solve for "t" in the continously compound formula

OpenStudy (anonymous):

isn't it 12ln3000?

OpenStudy (jdoe0001):

well... one sec

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

\(\bf A=Pe^{rt} \\ \quad \\ A=1,000,000\qquad P=3,000\qquad r=0.12 \\ \quad \\ 1,000,000=3,000e^{0.12t}\implies \cfrac{1,000}{3}=e^{0.12t} \\ \quad \\ ln\left(\frac{1,000}{3}\right)=0.12t\implies \cfrac{ln\left(\frac{1,000}{3}\right)=}{0.12}=t\)

OpenStudy (anonymous):

i knew i was missing a number i completely forgot about 1000000

OpenStudy (jdoe0001):

so you'll be a millionarie, but you'd be also about 70 years old too :P minus inflation you'd have less buying power anyway

OpenStudy (anonymous):

haha nice one! :) thank you for all your help i was stuck on it for an hour!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

one day i hope to know math as good as you it my weakest subject :/